# Advancing Security with Digital Twins: A Comprehensive Survey

### v1

@@ V-A An Introduction to Large Language Models
LLMs represent a groundbreaking advancement in AI, functioning as sophisticated deep learning systems capable of processing and generating human-like text. The evolution of these models began with the introduction of the ‘transformer’ architecture in 2017 by Google Brain, which marked a significant shift from traditional recurrent neural networks [114]. This revolutionary architecture processes text in parallel rather than sequentially, dramatically improving training efficiency and comprehension of lengthy text sequences. The core technological foundation of modern LLMs consists of transformer neural networks containing encoders and decoders with self-attention capabilities that extract meaning from text sequences and understand relationships between words and phrases. Unlike earlier models, transformers process entire sequences in parallel, allowing for GPU-accelerated training and reduced processing time.
Significant milestones in LLM development include Google’s bidirectional encoder representations from transformers (BERT) in 2018, which demonstrated unprecedented versatility through extensive training on natural language data [115]. OpenAI’s GPT series marked another crucial evolution phase, with GPT-3’s 2020 release representing a quantum leap in capabilities through massive scaling to 175 billion parameters. Their recent advances in GPT-4 and beyond have further enhanced LLMs through instruction tuning and conversational interfaces, making these technologies increasingly accessible through various platforms [116]. These models have demonstrated remarkable abilities in few-shot and zero-shot learning scenarios, setting new benchmarks in tasks ranging from text generation to semantic understanding.
LLMs have transformed numerous industries through their natural language processing capabilities. The versatility of LLMs is noticed in their ability to perform multiple language tasks - answering questions, summarizing documents, translating languages, and completing text sequences. This flexibility enables them to enhance traditional workflows across domains. Their remarkable ability to make predictions based on relatively small input prompts has enabled generative AI applications that produce human-like content in response to natural language instructions [117].
@@ V-B Integration of LLMs in DT Frameworks
Digital twins used across industries, from manufacturing to healthcare and urban planning, generate vast amounts of data that can be difficult to interpret and utilize effectively. LLMs address this challenge by bringing unprecedented levels of intelligence to these virtual environments. The integration of LLMs into these frameworks represents a significant advancement in the field, enhancing the capabilities of digital twins through natural language processing, reasoning, and data generation. This enables more intuitive interaction with DTs, automates the generation of scene description documents, and aids explainable decision-making across multiple domains. This substantially enhances the value proposition of DTs by converting data abundance into meaningful intelligence [118]. Figure 4 showcases the exploration of LLMs in DT systems so far, with the potential to go even beyond the following scopes:
@@ V-B1 Modeling and Simulation Enhancement
LLMs facilitate sophisticated descriptive, predictive, and prescriptive modeling by interpreting vast and complex datasets, thus holistically improving accuracy and reliability of predictions. TWIN-GPT, for example, is fine-tuned on a pre-trained LLM (ChatGPT) on clinical trial datasets to generate personalized digital twins for different patients [119]. This approach accounts for individual patient variations and disease complexities, producing data that closely aligns with diverse real-world scenarios, while also ensuring patient privacy. Additionally, they assist in system modeling and optimization through intelligent analysis of parameter interactions and system dynamics, and LLMs support efficient information retrieval and experimental design by automatically synthesizing insights from extensive simulations, streamlining hypothesis testing and data-driven decision-making in complex simulation scenarios. For instance, a recent study proposes using LLM agents within a shopping mall twin to simulate human behaviors and thermal preferences [120]. These LLM agents interact with the digital twin environment to provide realistic data for optimizing HVAC systems.
@@ V-B2 Natural Language Interaction and Automation
LLMs bring natural language understanding to DT systems, greatly improving human-system interaction and automation. Natural language query interfaces enable intuitive interactions, allowing users to communicate with digital twins without requiring deep technical expertise. LLMs can automate the generation of scene description documents for DTs, reducing the need for manual processes and domain expertise. ChatTwin, a system that utilizes GPT-4 to automate the generation of scene description documents for data center DTs, employs a segment-and-generate workflow to address the challenges of long-text generation, achieving a higher success rate and completeness compared to baseline prompting methods [121]. This directly demonstrates the use of LLMs in automatically creating digital twin specifications. LLMs can also provide an explainability platform, generating natural language explanations of the system’s decision-making by leveraging domain-specific knowledge bases [58]. Retrieval-augmented generation allows LLMs to utilize technical documents, scientific papers, and computer code to explain the decisions being taken by the system in real-time. LLMs further facilitate effective communication and automate complex analytical processes, bridging the gap between technical complexity and end-user accessibility.
@@ V-B3 Semantic Communication and Computation
Semantic communication within digital twin frameworks is greatly improved by integrating LLMs, notably through the concept of digital twin semantic networks, which enable more efficient communication and computation within DT environments [122]. This approach reshapes both intra-twin and inter-twin communication architectures through semantic-level frameworks that unify efficient communication and computation. By enabling DTs to communicate and compute at the semantic level rather than through raw data exchange, this framework significantly reduces bandwidth requirements and computational overhead.
@@ V-B4 Resource-Constrained Deployment Strategies
Resource-constrained environments pose unique challenges for deployment, which can be effectively addressed through integration with LLMs. This is addressed via a mini-giant LLM collaboration scheme designed for resource-constrained environments. This approach involves edge data fine-tuning and instruction prompts, where larger models provide rich knowledge information while smaller models handle personalized expertise updating. This collaboration enhances the capabilities of DTs for multimodal data processing, scalability, and adaptability without overwhelming computational resources. Moreover, lightweight model updating techniques facilitated by LLMs enable continuous and efficient model refinement, ensuring up-to-date system intelligence without extensive computational resources.
@@ V-B5 Security
Security is critically enhanced by incorporating large language models into digital twin environments. LLM-powered tools facilitate automatic generation of secure hardware description code, improving accuracy and security at the hardware design level. Advanced property extraction mechanisms, such as neural security property generation, utilize the semantic understanding capabilities of LLMs to reliably identify and extract hardware security properties. Furthermore, LLM-based assertion generation tools support proactive hardware vulnerability detection, while debugging and syntax error correction solutions enable faster resolution of security-critical errors, strengthening the resilience of digital twin systems. Recent research has extensively investigated employing LLMs to enhance hardware security assurance. Thakur et al. demonstrated how LLMs, such as CodeGen, can automate synthesizable Verilog code generation [123]. Their comparative study, involving fine-tuning models against a large-scale Verilog corpus, revealed competitive performance against established models like ChatGPT, showcasing the feasibility of automating hardware digital twin creation at the register transfer level. Kande et al. investigated OpenAI’s code-davinci-002 model for automatic generation of hardware security assertions [124]. Their research explored how varying prompt detail levels impacted the quality of generated SystemVerilog assertions, essential components in hardware verification, which directly influence digital twin security assurance. Advancing the role of LLMs in hardware debugging, Tsai et al. introduced RTLFixer, a framework employing LLMs to automatically identify and rectify syntax errors in RTL code [125]. Such automated correction tools substantially streamline the hardware design and validation processes for digital twins, significantly reducing manual debugging overhead and enhancing overall system security.
@@ V-C Future Prospects of LLMs and DTs in Security
Despite promising advancements, integrating large language models with digital twins introduces several critical challenges that warrant ongoing attention. Among these are security concerns arising from semantic-level operations, potentially opening new vulnerabilities and attack vectors unique to LLM-driven digital twin frameworks. Moreover, computational efficiency remains a pivotal issue, particularly in resource-constrained deployment scenarios where computing resources and energy budgets are limited. Ensuring the accuracy, reliability, and robustness of insights and security assurances derived from LLM-generated outputs requires robust validation frameworks and rigorous testing methodologies, especially within safety-critical domains such as healthcare, manufacturing, and infrastructure management.
The integration of LLMs with DTs represents a promising frontier for advancing security measures across various cyber-physical systems. Utilizing the logical reasoning and generative capabilities of LLMs, future digital twin frameworks can potentially offer unprecedented levels of security analysis, anomaly detection, and automated response to threats. Specialized models trained explicitly for security-related applications could significantly enhance the accuracy and efficiency of threat identification, vulnerability assessment, and proactive defense mechanisms. Emerging research directions include developing multimodal LLMs capable of interpreting and analyzing combined textual, numerical, and visual data to provide comprehensive, real-time security evaluations of digital twin environments. Furthermore, the adoption of federated learning methods holds promise for enabling decentralized and privacy-preserving intelligence across networks of interconnected twins, thereby improving resilience against centralized points of failure and coordinated cyberattacks. We anticipate increased convergence between physical systems and their digital representations, augmented by sophisticated security-aware language models. This evolution will not only enhance situational awareness and rapid incident response but also facilitate proactive risk mitigation strategies, thus significantly strengthening the security posture of critical infrastructure and cyber-physical systems.
@@ VI-B Data Acquisition
DTs rely heavily on high-quality, real-time data streams to accurately mirror the behavior and condition of their physical counterparts. However, missing data points, sensor noise, or inconsistencies in measurements can significantly degrade the accuracy and reliability of the twins’ predictions and decision-making capabilities. As a result, substantial effort must be invested into data cleaning, validation, and preprocessing to ensure the digital twin operates as intended.
In the semiconductor industry, this challenge is further compounded by data accessibility issues, as key stakeholders such as foundries, testing facilities, and equipment vendors may be reluctant to share proprietary or sensitive operational data due to intellectual property concerns, data privacy, or competitive reasons. This reluctance can hinder the construction of comprehensive digital twins by creating blind spots in critical process stages or component behavior. Without full visibility into the system, the digital twin may be forced to rely on approximations or incomplete models, reducing its value in optimization and root-cause analysis.
@@ VI-C Security and Privacy
VI-D Token Size and Hallucination Problem
Large language models are inherently constrained by a fixed token window, which limits the amount of input data they can process in a single inference. This limitation poses a significant challenge when analyzing digital twins of complex hardware systems, as large-scale system logs, event traces, or multi-modal telemetry data often exceed this limit. When truncated inputs omit critical contextual information, the model’s reasoning becomes less robust, increasing the risk of generating incomplete or misleading outputs. This is particularly problematic in security-sensitive applications, where overlooking a subtle dependency or omitting a causal signal can result in false negatives or misdiagnosed vulnerabilities.
In addition to context loss, LLMs are susceptible to hallucination, generating information that is syntactically plausible but semantically incorrect. For example, when tasked with identifying side-channel vulnerabilities in a hardware digital twin, an LLM may fabricate data dependencies or invent leakage channels that do not exist in the actual system. Such hallucinations not only reduce analytical trustworthiness but may also lead to wasted debugging cycles, misallocated resources, or worse, neglect of real threats.
In order to mitigate these limitations, techniques such as Retrieval-Augmented Generation (RAG)[126] and extended context models like Longformer[127] can be used to integrate longer or dynamically indexed context windows. Additionally, domain-specific fine-tuning [128] ensures that the LLM internalizes accurate patterns from hardware-centric corpora. Human-in-the-loop strategies [129] provide a final layer of validation, enabling real-time correction of hallucinated outputs and alignment with operational ground truth. Together, these interventions help improve the semantic fidelity, robustness, and reliability of LLMs deployed within workflows.
@@ VII Research Trends
These milestones are not merely historical footnotes; they reflect the maturation of digital twins from an abstract concept into a multidimensional technology with real-world impact. Each turning point — whether the entry into manufacturing, healthcare, or security — has incrementally extended the capabilities and scope of DT systems. As the fusion of digital twins with LLMs becomes more prominent, we are likely to witness the emergence of intelligent, adaptive DTs capable of semantic reasoning, self-optimization, and autonomous decision-making [119, 121]. The convergence of these technologies represents a pivotal shift, potentially transforming DTs from passive monitoring systems into active, explainable, and secure agents across critical domains.

### v2

@@ V-A An Introduction to Large Language Models
LLMs represent a groundbreaking advancement in AI, functioning as sophisticated deep learning systems capable of processing and generating human-like text. The evolution of these models began with the introduction of the ‘transformer’ architecture in 2017, which marked a significant shift from traditional recurrent neural networks [108]. Another significant milestone in LLM development is the introduction of bidirectional encoder representations from transformers (BERT) in 2018, which demonstrated unprecedented versatility through extensive training on natural language data [109]. OpenAI’s GPT series marked another crucial evolution phase, with GPT-3’s 2020 release followed by GPT-4 more recently, has further enhanced LLMs through instruction tuning and user-friendly conversational interfaces [110]. LLMs have transformed numerous industries through their natural language processing capabilities, performing multiple language tasks - answering questions, summarizing documents, translating languages, and completing text sequences. This flexibility enables them to enhance traditional workflows across domains. Their remarkable ability to make predictions based on relatively small input prompts has enabled generative AI applications that produce human-like content in response to natural language instructions [111].
@@ V-B Integration of LLMs in DT Frameworks
From manufacturing to healthcare and urban planning, DTs generate vast amounts of data that can be difficult to interpret and utilize effectively. Integrating LLMs into these frameworks allows more intuitive interaction with DTs and augments explainable decision-making across multiple domains. Figure 4 showcases the exploration of LLMs in DT systems so far, with the potential to go even beyond the following scopes:
@@ V-B1 Modeling and Simulation Enhancement
LLMs can conduct descriptive, predictive, and prescriptive modeling by interpreting vast and complex datasets, thus holistically improving the accuracy and reliability of predictions. TWIN-GPT, for example, is fine-tuned on a pre-trained LLM on clinical trial datasets to generate personalized digital twins for different patients [112]. LLMs also assist in system modeling and optimization through intelligent analysis of parameter interactions and system dynamics, and LLMs support information retrieval and experimental design by automatically synthesizing insights from extensive simulations, streamlining hypothesis testing and data-driven decision-making in complex simulation scenarios. For instance, a recent study proposes using LLM agents within a shopping mall twin to simulate human behaviors and thermal preferences [113]. These LLM agents interact with the DT environment for optimizing HVAC systems.
@@ V-B2 Natural Language Interaction and Automation
LLMs bring natural language understanding to DT systems and improve human-system interaction and automation. Natural language query interfaces allow users to communicate with digital twins without requiring deep technical expertise. LLMs can automate the generation of scene description documents for DTs, reducing the need for manual processes and domain expertise. ChatTwin, a conversational system that automates the generation of scene description documents for data center DTs, showcases the ability of LLMs to effectively model and generate data center scenes of the DT [114]. LLMs can also provide an explainability platform, generating natural language explanations of the system’s decision-making by utilizing domain-specific knowledge bases. Retrieval-augmented generation (RAG) allows LLMs to utilize technical documents, scientific papers, and code to explain real-time decisions taken by the system.
@@ V-B3 Semantic Communication and Computation
Semantic communication within digital twin frameworks is improved by integrating LLMs, notably through the concept of digital twin semantic networks by more efficient communication and computation within DT environments [115]. This approach reshapes both intra-twin and inter-twin communication architectures through semantic-level frameworks that unify efficient communication and computation. By enabling DTs to communicate and compute at the semantic level rather than through raw data exchange, this framework significantly reduces bandwidth requirements and computational overhead.
@@ V-B4 Resource-Constrained Deployment Strategies
Resource-constrained environments pose unique challenges for deployment, which can be effectively addressed through integration with LLMs. A mini-giant LLM collaboration scheme designed for resource-constrained environments involves edge data fine-tuning and instruction prompts, where larger models provide rich knowledge information while smaller models handle personalized expertise updating. Moreover, lightweight model updating techniques by LLMs enable continuous and efficient model refinement, ensuring up-to-date system intelligence without extensive computational resources.
@@ V-B5 Security
Security is critically enhanced by incorporating LLMs into DT environments. LLMs can automatically generate secure hardware description code, improving accuracy and security at the hardware design level. Thakur et al. demonstrated how LLMs, such as CodeGen, can automate synthesizable Verilog code generation [116] at the register transfer level (RTL). Advanced property extraction mechanisms, such as neural security property generation, utilize the semantic understanding capabilities of LLMs to reliably identify and extract hardware security properties. Furthermore, LLM-based assertion generation tools support proactive hardware vulnerability detection, while debugging and syntax error correction solutions enable faster resolution of security-critical errors. Kande et al. investigated LLMs for automatic generation of hardware security assertions [117]. Their research explored the use of natural language prompts in LLMs to produce SystemVerilog assertions, which are indeed challenging to write manually. Tsai et al. introduced RTLFixer, a framework integrating LLMs to automatically identify and rectify syntax errors in RTL code [118]. This framework also notably utilizes RAG and an advanced prompting technique that results in exceptional proficiency in resolving syntax errors in code.
@@ V-C Future Prospects of LLMs and DTs in Security
Despite such promising advancements, integrating LLMs with DTs has introduced several critical issues along the way. Notably, in resource-constrained deployment scenarios where compute is limited, computational efficiency remains an area of concern. Another challenge is the accuracy and reliability of the LLM-generated outputs that require robust validation frameworks, especially within safety-critical domains like healthcare and manufacturing.
As these challenges are gradually acknowledged and dealt with, LLMs and DTs together truly represent a strong unified solution for advancing security measures across cyber-physical systems. With the logical reasoning and natural language understanding capabilities of LLMs, the DT frameworks in the future can offer improved security analysis and anomaly detection. Recent trends include training specialized models for more accurate security-related applications and exploring the multimodality of LLMs to interpret both textual and visual data to provide real-time security evaluations of DT environments. We anticipate that this convergence between LLMs and DTs will continue to significantly strengthen the security of critical infrastructure.
@@ VI-B Data Acquisition
DTs rely heavily on high-quality, real-time data streams to accurately mirror the behavior and condition of their physical counterparts. However, missing data points, sensor noise, or inconsistencies in measurements can significantly degrade the accuracy and reliability of DT’s predictions and decision-making capabilities. As a result, substantial effort must be invested into data cleaning, validation, and preprocessing to ensure that the DT operates as intended. This challenge is further compounded in the semiconductor industry due to data accessibility issues, as key stakeholders such as foundries, testing facilities, and equipment vendors may be reluctant to share proprietary or sensitive operational data due to intellectual property concerns, data privacy, or competitive reasons. This reluctance can hinder the construction of comprehensive DTs by creating blind spots in critical process stages or component behavior. Without full visibility into the system, the DT may be forced to rely on approximations or incomplete models, reducing its value in optimization and root-cause analysis.
@@ VI-C Security and Privacy
VI-D Token Size and Hallucination Problems of LLMs
While LLMs are considered to have the relatively untapped potential to enhance the security of DTs, they are inherently constrained by a fixed token window that limits the amount of input data they can process in a single inference. This poses a significant challenge while analyzing DTs of complex hardware systems because large-scale system logs, event traces, or multi-modal telemetry data often exceed this limit. In such situations, the inputs get truncated and critical context gets omitted in the process. In security-sensitive applications, this might lead to generating incomplete or even misleading outputs.
Moreover, as probabilistic token predictors, LLMs are prone to generating information that is syntactically plausible but semantically incorrect. This is termed as hallucination. For example, when tasked with identifying side-channel vulnerabilities in a hardware digital twin, an LLM may hallucinate and fabricate leakage channels that do not even exist in the actual system. While this feature comes inherently with the non-deterministic nature of LLMs, it significantly reduces the reliability of the generated responses, especially in security.
@@ VII Research Trends
These milestones are not merely historical footnotes; they reflect the maturation of digital twins from an abstract concept into a multidimensional technology with real-world impact. Each turning point — whether the entry into manufacturing, healthcare, or security — has incrementally extended the capabilities and scope of DT systems. As the fusion of digital twins with LLMs becomes more prominent, we are likely to witness the emergence of intelligent, adaptive DTs capable of semantic reasoning, self-optimization, and autonomous decision-making. The convergence of these technologies represents a pivotal shift, potentially transforming DTs from passive monitoring systems into active, explainable, and secure agents across critical domains.